The length of a rectangle is 30cm less than three times the width. The perimeter of this rectangle is 180cm. What are the dimensions of this rectangle?
Answer by checkley77(12844)

L=3W-30
2L+2W=30
2(3W-30)+2W=30
6W-60+2W=30
8W=30+60
8W=90
W=90/8
W=11.25 CM. IS THE WIDTH.
L=3*11.25-30=33.75-30=3.75 CM IS THE LENGTH.
PROOF:
2*3.75+2*11.25=30
7.5+22.5=30
30=30